# MHRC-Bench: A Multilingual Hardware Repository-Level Code Completion benchmark

### v1

@@ Abstract
Large language models (LLMs) have achieved strong performance on code completion tasks in general-purpose programming languages. However, existing repository-level code completion benchmarks focus almost exclusively on software code and largely overlook hardware description languages. In this work, we present MHRC-Bench, consisting of MHRC-Bench-Train and MHRC-Bench-Eval, the first benchmark designed for multilingual hardware code completion at the repository level. Our benchmark targets completion tasks and covers three major hardware design coding styles. Each completion target is annotated with code-structure-level and hardware-oriented semantic labels derived from concrete syntax tree analysis. We conduct a comprehensive evaluation of models on MHRC-Bench-Eval. Comprehensive evaluation results and analysis demonstrate the effectiveness of MHRC-Bench.
MHRC-Bench: A Multilingual Hardware Repository-Level Code Completion benchmark
Qingyun Zou1, Jiahao Cui1, Nuo Chen1, Bingsheng He1, Weng-Fai Wong1 1National University of Singapore qingyunzou@u.nus.edu
Figure 1: Average EM and ES comparison of different LLMs on the MHRC-Bench-Eval benchmark.
Figure 2: Overview of our hardware code completion benchmark across multiple hardware description paradigms. We present three representative examples from generator-based hardware construction languages (e.g., Chisel), high-level synthesis (HLS) languages (e.g., Xilinx HLS), and Register-Transfer Level (RTL) design languages (SystemVerilog). For each example, both in-file and cross-file contexts are provided, with the completion target explicitly marked. Models are required to predict the correct completion given the provided contexts, where “<TARGET>” indicates the completion position. More examples are introduced in Appendix A.10
@@ 1 Introduction
Recent advancements in large language models (LLMs) have led to significant progress in code-related tasks such as code understanding, generation, and completion (Liu et al., 2023; Nam et al., 2024; Liu et al., 2024a; Pan et al., 2024; Raihan et al., 2025; Yang et al., 2024).
These models, trained on vast datasets encompassing both natural language and source code, have demonstrated remarkable proficiency in modeling programming languages as structured sequences and predicting logical code continuations (Liang et al., 2024; Li et al., 2025; Guo et al., 2024; Hui et al., 2024; Liu et al., 2024b; Abedu et al., 2025; Hua et al., 2025). Consequently, LLM-based code completion has emerged as a valuable tool in real-world software development environments.
More recently, research efforts have shifted toward repository-level code completion, where models are required to leverage cross-file information such as imports, symbol definitions, and project structure. Wang et al. (2024) train a dedicated retriever to identify and retrieve relevant code fragments that can be incorporated into LLM-based code completion. In a complementary direction, Deng et al. (2025) equip LLMs with external tool APIs to more effectively utilize repository-level context during the code completion process. While recent work has explored Verilog code completion (Wu et al., 2025b; Allam and Shalan, 2024), existing benchmarks remain narrowly focused on a single hardware language. Inspired by Liu et al. (2025b), we propose MHRC-Bench, the first multilingual repository-level benchmark for hardware code completion, consisting of MHRC-Bench-Train and MHRC-Bench-Eval.
Using the MHRC-Bench-Train dataset to fine-tune code LLMs, we observe that billion-scale fine-tuned models such as Qwen2.5-Coder-7B achieve substantially higher EM and ES scores than their pretrained counterparts, whose EM is nearly zero on hardware code completion tasks. After fine-tuning, these models also outperform larger general-purpose models, including GPT-5, Grok-4, Gemini 2.5 Pro, and DeepSeek V3.2, as shown in Figure 1. This demonstrates that MHRC-Bench-Train effectively enhances the hardware code completion capabilities of smaller-scale code LLMs.
@@
The benchmark and related code are publicly available. The contributions are summarized as follows:
@@
  • We construct MHRC-Bench-Train, a training dataset designed to enhance repository-level hardware code completion. Fine-tuning on MHRC-Bench-Train leads to substantial performance gains, even enabling smaller open-weight models to outperform strong commercial baselines.
  • We conduct comprehensive evaluations across a wide range of models, including commercial LLMs and open-source code LLMs, as well as state-of-the-art retrieval-based methods. Our analysis reveals key performance trends and insights into LLM behavior in hardware code completion.
2 Related Works and Background
@@
Recent repository-level code completion methods (Deng et al., 2025; Li et al., 2025; Liu et al., 2024b; Liang et al., 2024; Liu et al., 2025d; Wu et al., 2025b; Wang et al., 2024; Zhang et al., 2025) aim to retrieve related code snippets across files within a repository, where existing datasets mainly focus on software engineering task. For example, M2RC-Eval (Liu et al., 2025b) supports Java, Python and so on these related software codes completion.
2.2 LLM for EDA.
Recent advancements in applying large language models (LLMs) to electronic design automation (EDA) tasks, such as RTL/HLS code generation and hardware verification, have shown significant promise. Notable studies in NLP venues include SynC-LLM (Liu et al., 2025c), which uses diffusion models and LLM prompting to generate realistic digital circuits, improving downstream EDA model performance. Wu et al. (2025a) introduces a multi-agent LLM system that enhances the reliability of multi-step EDA tool flows. DeepRTL2 (Liu et al., 2025e) showcases LLMs’ versatility in RTL generation and embedding-based tasks, setting new benchmarks. Lastly, AssertionBench (Pulavarthi et al., 2025) presents a benchmark for Verilog designs and assertion generation, highlighting progress and areas for improvement. These studies demonstrate the growing impact of LLMs in automating hardware design. However, these works only focus on kernel-level code rather than repository-level codes and only concentrate on one hardware design language.
@@ 3.2 Dataset Comparison
  • High-Level Synthesis (HLS) languages. They describe hardware behavior using C/C++ like abstractions augmented with synthesis directives such as Xilinx HLS C/C++ and SystemC;
  • Generator-based hardware construction languages. They enable parameterized and programmable generation of RTL designs such as Chisel.
Compared with existing hardware repository-level benchmark-RTL-Repo Allam and Shalan (2024), MHRC-Bench enables a comprehensive evaluation of LLMs across different abstraction levels, programming styles, and domain-specific constraints in hardware-oriented code completion.
Table 1: Comparison with repository-level hardware code datasets. Lang indicates languages, FG indicates fine-grained annotation.
Benchmark Lang FG Target Samples
RTL-Repo RTL × Single-line 4098
Ours RTL, HLS, Generator Single & Multi-line 47,175
@@
Syntactically complete. Prior work Ding et al. (2023) often selects random character spans as completion targets, which may break syntactic integrity. In contrast, recent studies (Liu et al., 2025b) show that developers expect LLMs to generate syntactically complete code units (e.g., full lines or blocks). Following this insight, MHRC-Bench-Eval uses tree-sitter [^1] https://github.com/tree-sitter/tree-sitter to parse each file into a concrete syntax tree and randomly selects a syntactic node as the completion target, using its exact source span as the ground-truth output.
Data Independence and Inference difficulty. To ensure data independence and inference difficulty: (1) The repositories and files used for the training, validation, and test sets are mutually exclusive, with no overlap. (2) We ensure that 40% of completion targets are between 2 and 5 lines inclusive. (3) The target code should not exclusively include whitespace, debug message and comments. (4) For each file, we select exactly one target position.
@@ 3.4 Dataset Statistics
Table 2: Dataset statistics of MHRC-Bench. Val means validation. Cross-file means cross-file dependencies. Every column lists its numbers of completion targets.
Language Repos Train Val Test Cross-file
Chisel 48 2,573 119 319 3.12
Verilog/SV 247 25,586 107 492 4.02
VHDL 239 13,039 104 491 4.34
HLS 50 3,766 123 456 5.85
@@ 3.5 Dataset Categories
To better analyze LLM’s performance in a fine-grained manner, we further provide two types of fine-grained annotations include code structural level and hardware semantic level) for each completion target as shown in Figure 3.
@@
Hardware semantic level. We categorize the target code regions according to the hardware design and verification flow in Table 3, which broadly consists of design coding and testbench-based verification. Unlike software programs, where execution is inherently sequential, hardware code describes a parallel architecture or circuit with explicit structural, temporal, and interface semantics. This fundamental difference motivates a hardware-specific categorization of completion targets.
For design coding, we divide the target code into the following categories: Design Structure, Declaration and Definition, Storage Block, Computation Block, Control Flow Block, and Interface. These categories capture the key aspects of hardware design, ranging from structural hierarchy and state storage to data-path computation, control logic, communication interfaces.
In addition, we explicitly model verification code, which plays a critical role in hardware development but is largely absent in software benchmarks. We further decompose the verification code into several sub-types, including Property and Assertion Specification, Testbench Stimulus and Environment Modeling, and Monitoring and Checking Logic. These sub-types collectively describe how correctness properties are specified and how its behavior is observed and validated during simulation or formal verification. The distribution of hardware semantic categories is presented in Appendix A.2.
@@ 4.1 Evaluation Models and Metrics
Code model. We evaluate Code LLMs of different scale. The evaluated models include DeepSeek-Coder-MQA-5.7B Guo et al. (2024), DeepSeek-Coder-base-6.7B Guo et al. (2024), Qwen2.5-Coder-3B-base Hui et al. (2024), Qwen2.5-Coder-7B-base Hui et al. (2024) and Qwen2.5-Coder-14B-base Hui et al. (2024).
Common model. We evaluate commerical models including GPT-5 Leon (2025), Gemini 2.5 Pro Comanici et al. (2025), Grok4 Adinath and IS (2025), and Deepseek V3.2 Liu et al. (2025a).
Metrics. We compare the generated code with the reference and compute the exact match (EM), edit similarity (ES) and Bilingual Evaluation Understudy (BLEU) following Liu et al. (2025b); Zhang et al. (2023). Their detailed implementations are defined in A.3
Model Chisel V/SV VHDL HLS
EM ES BLEU EM ES BLEU EM ES BLEU EM ES BLEU
Qwen2.5-Coder-3B 0.0 7.9 5.4 0.2 7.9 2.8 0.0 6.6 2.9 0.0 5.5 4.1
+Tuning 24.8 54.3 50.6 34.8 65.5 56.5 28.2 59.1 60.2 33.1 55.0 55.7
Qwen2.5-Coder-7B 0.0 14.8 4.9 0.0 12.7 3.3 0.0 13.0 3.8 0.2 14.1 4.5
+Tuning 26.6 66.7 56.6 39.0 70.8 60.1 34.5 71.9 64.9 35.5 69.5 58.3
Qwen2.5-Coder-14B 0.0 16.5 6.9 1.0 18.5 7.5 0.4 15.8 5.7 1.1 16.0 6.3
+Tuning 31.3 68.9 59.6 41.9 74.4 64.2 35.3 73.8 67.7 38.2 71.5 61.3
DS-Coder-5.7B 0.0 13.2 3.7 0.4 8.8 2.6 0.0 9.0 3.2 0.0 9.7 2.8
+Tuning 24.8 61.2 50.1 37.6 71.7 61.0 30.7 71.1 63.0 33.1 67.7 55.9
DS-Coder-6.7B 0.0 5.4 3.4 0.2 5.7 2.5 0.0 5.6 2.6 0.0 3.6 2.2
+Tuning 25.7 56.5 52.6 37.8 67.4 61.2 31.5 61.1 63.9 36.8 60.4 58.2
Table 4: Exact Match (EM %), Edit Similarity (ES %), and BLEU scores % across different hardware description languages after fine-tuning for code LLMs.
Model Retrieval Method Chisel V/SV VHDL HLS
EM ES BLEU EM ES BLEU EM ES BLEU EM ES BLEU
GPT-5 No RAG 24.8 56.7 45.9 34.5 60.9 49.5 15.4 52.0 41.7 31.4 63.7 55.3
RepoCoder 27.6 59.6 49.9 35.4 61.3 50.2 16.7 54.1 43.5 34.4 63.9 55.5
GraphCoder 13.9 44.2 36.4 15.7 40.1 31.8 7.5 35.7 29.0 24.9 49.5 41.1
RLCoder 27.8 60.2 50.3 36.2 61.7 51.4 17.4 55.3 45.7 36.9 72.9 60.1
Gemini 2.5 Pro No RAG 33.2 59.9 49.8 33.2 59.9 49.8 16.3 50.8 41.4 34.6 63.8 57.3
RepoCoder 33.7 61.7 54.7 34.8 59.6 50.5 18.0 52.3 42.3 39.4 65.3 59.1
GraphCoder 13.3 50.8 44.0 22.2 50.5 43.2 9.1 45.0 39.8 29.1 57.8 50.8
RLCoder 34.6 63.2 55.1 35.3 60.2 51.3 19.2 53.7 43.5 41.2 66.9 60.2
Grok4 No RAG 22.4 53.7 42.1 31.4 56.8 46.7 14.3 47.8 36.3 31.6 60.5 52.3
RepoCoder 23.8 54.7 44.3 32.0 56.1 45.5 15.7 49.3 37.4 32.3 61.1 53.7
GraphCoder 12.7 35.4 30.5 16.5 38.1 31.8 8.1 35.3 30.3 18.0 38.8 33.1
RLCoder 25.6 55.2 43.1 33.4 57.2 47.9 16.8 51.2 38.2 33.3 62.2 61.3
Deepseek V3.2 No RAG 22.6 51.2 39.8 28.7 54.1 42.2 13.3 43.0 30.0 30.3 57.3 49.5
RepoCoder 23.2 52.1 41.6 28.3 52.6 40.7 14.2 42.8 29.2 33.3 57.9 50.8
GraphCoder 9.3 39.5 32.8 15.1 39.1 32.1 9.9 39.1 33.0 22.7 47.9 40.3
RLCoder 24.3 53.1 40.2 29.4 45.1 42.4 14.2 43.6 31.3 34.3 59.2 51.0
Table 5: EM %, ES %, and BLEU scores % with different retrieval methods across hardware description languages.
@@ 4.2 Experimental Setup
(a) Impact of dataset size on EM performance.
(b) Impact of dataset size on ES performance.
(c) Model performance and the number of lines to be completed.
Figure 4: Impact of training data scale and completion length on hardware code completion.
(a) EM of GPT-5 without retrieval augmentation (No-RAG) under increasing prompt lengths.
(b) EM of post-trained Qwen2.5-Coder-7B under increasing prompt lengths.
(c) Model performance and code structure depth.
Figure 5: Effects of context length and code structure depth on hardware code completion.
(a) EM performance of different categories.
(b) ES performance of different categories.
Figure 6: Performance of GPT-5 and post-trained Qwen2.5-Coder-7B across hardware semantic categories.
@@
Baseline. We provide the model only with the file that contains the completion cursor. Without explicit inter-file context, the model must depend on its pretrained ability to produce the completion.
Tuning. To further improve the performance of repository-level code completion, we fine-tune code LLMs on training dataset supervised fine-tuning (SFT) with detailed configuration in Appendix A.3.
@@
Retrieval Framework. There are many different repository-level code completion methods including RepoCoder (Zhang et al., 2023), GraphCoder (Liu et al., 2024b), and RLCoder Wang et al. (2024). We evaluate these frameworks with different models on MHRC-Bench-Eval. The total input window is fixed at 2,048 tokens for retrieved context and local context for all the models.

### v2

@@ Abstract
Large language models (LLMs) have achieved strong performance on code completion tasks in general-purpose programming languages. However, existing repository-level code completion benchmarks focus almost exclusively on software code and largely overlook hardware description languages. In this work, we present MHRC-Bench, consisting of MHRC-Bench-Train and MHRC-Bench-Eval, the first benchmark designed for multilingual hardware code completion at the repository level. Our benchmark targets completion tasks and covers three major hardware design coding styles. Each completion target is annotated with code-structure-level and hardware-oriented semantic labels derived from concrete syntax tree analysis. We conduct a comprehensive evaluation of models on MHRC-Bench-Eval. Our evaluation reveals insights into the differing behaviors of LLMs in hardware code completion across multiple hardware languages.
Machine Learning, ICML
Figure 1: Overview of our hardware code completion benchmark across multiple hardware description paradigms. We present three representative examples from domain-specific language (DSL) (e.g., Chisel), high-level synthesis (HLS) languages (e.g., Xilinx HLS), and Register-Transfer Level (RTL) design languages (SystemVerilog). For each example, both in-file and cross-file contexts are provided, with the completion target explicitly marked. Models are required to predict the correct completion given the provided contexts, where “<TARGET>” indicates the completion position. More examples are introduced in Appendix A.11
@@ 1 Introduction
Recent advancements in large language models (LLMs) have led to significant progress in code-related tasks such as code understanding, generation, and completion (Liu et al., 2023b; Nam et al., 2024; Liu et al., 2024a; Pan et al., 2024; Raihan et al., 2025; Yang et al., 2024).
These models, trained on vast datasets encompassing both natural language and source code, have demonstrated remarkable proficiency in modeling programming languages as structured sequences and predicting logical code continuations (Liang et al., 2024; Li et al., 2025; Guo et al., 2024; Hui et al., 2024; Liu et al., 2024c; Abedu et al., 2025; Hua et al., 2025). Consequently, LLM-based code completion has emerged as a valuable tool in real-world software development environments.
More recently, research efforts have shifted toward repository-level code completion, where models are required to leverage cross-file information such as imports, symbol definitions, and project structure. Wang et al. (2024) train a dedicated retriever to identify and retrieve relevant code fragments that can be incorporated into LLM-based code completion. In a complementary direction, Deng et al. (2025) equip LLMs with external tool APIs to more effectively utilize repository-level context during the code completion process. Despite their success in software code, LLMs’ performance on hardware code remains unclear, as it differs fundamentally in execution semantics, structural parallelism, and cross-file dependency patterns. In particular, hardware code completion requires understanding static structure, signal connectivity, and design intent rather than purely sequential execution. We introduce MHRC-Bench including MHRC-Bench-Train and MHRC-Bench-Eval, a multilingual repository-level benchmark designed to expose structural and semantic failure modes of LLMs in hardware code completion.
Using the MHRC-Bench-Train dataset to fine-tune code LLMs, we observe that billion-scale fine-tuned models such as Qwen2.5-Coder-7B achieve substantially higher EM and ES scores than their pretrained counterparts, whose EM is nearly zero on hardware code completion tasks. After fine-tuning, the post-trained Qwen2.5-Coder-7B also outperform larger general-purpose models such as GPT-5 about EM, ES, and compilation pass rate. with the same input window. This demonstrates that MHRC-Bench-Train effectively enhances the hardware code completion capabilities of smaller-scale code LLMs.
@@
The benchmark, related code and fine-tuned models are publicly available. The contributions are summarized as follows:
@@
  • We construct MHRC-Bench-Train, a training dataset designed to enhance repository-level hardware code completion. Fine-tuning on MHRC-Bench-Train leads to substantial performance gains, even enabling smaller open-weight models to outperform larger general models.
  • We conduct comprehensive evaluations across a wide range of models, including general LLMs and open-source code LLMs, as well as state-of-the-art retrieval-based methods. Our analysis reveals key performance trends and insights into LLM behavior in hardware code completion.
Table 1: Comparison with hardware code benchmark. FG indicates fine-grained annotation. Train means if it contains training dataset. NL means natural language. RTL means register transfer level. HLS means High-Level Synthesis. DSL means domain-specific language.
Benchmark Level Task Type Input Languages FG Train Output Data types Samples
VerilogEval Module Code Gen NL Spec RTL × Module Code Synthetic and Real-world 8,502
RTLLM Module Code Gen NL Spec RTL × × Module Code Real-world 30
HLS-Eval Module Code Gen NL Spec or C Code HLS × × Module Code Real-world 94
HDLEval Module Code Gen NL Spec RTL, DSL × × Module Code Real-world 461
RTL-Repo Repository Code Completion Repository Code RTL × Single-line Real-world 4098
Ours Repository Code Completion Repository Code RTL,HLS,DSL Single and Multi-line Real-world 47,175
2 Related Works
@@
Recent repository-level code completion methods (Deng et al., 2025; Li et al., 2025; Liu et al., 2024c; Liang et al., 2024; Liu et al., 2025c; Wang et al., 2024; Zhang et al., 2025; Liu et al., 2025b) aim to retrieve relevant code snippets across multiple files within a repository. However, existing benchmarks and datasets predominantly focus on software engineering tasks, while largely overlooking hardware-oriented code completion scenarios.
2.2 Benchmarking LLMs for Hardware Code.
Existing work on benchmarking LLMs for hardware design code generation has primarily focused on isolated, module-level synthesis tasks. VerilogEval (Liu et al., 2023a) evaluates models on short Verilog problems using simulation-based correctness checks, while RTLLM (Lu et al., 2024) introduces natural-language-to-RTL tasks evaluated with pass@k metrics over handcrafted examples. HLS-Eval (Abi-Karam and Hao, 2025) assesses LLMs on HLS code generation from C code and textual specifications. More recent efforts such as HDLEval (Kashanaki et al., 2024) broaden the scope to include multiple HDL dialects and language-agnostic prompts, but still focus on specification-to-module generation under constrained conditions. Repository-level benchmarks like RTL-Repo (Allam and Shalan, 2024) remain limited to a single HDL, typically evaluate only one-line completions, and lack fine-grained analysis of LLM performance.
@@ 3.2 Dataset Comparison
  • High-Level Synthesis (HLS) languages. These describe hardware behavior using C/C++-like abstractions, often augmented with synthesis directives. Among them, Xilinx HLS C/C++ is widely used, so we include it in our benchmark.
  • Domain-Specific Languages (DSLs). DSLs enable parameterized and programmable generation of RTL designs. Chisel is a representative example, and we include it in our benchmark for its growing adoption.
As shown in Table 1, MHRC-Bench contains the largest number of data samples, supports multilingual repository-level hardware code, and provides fine-grained annotations, enabling comprehensive evaluation of LLMs in hardware code completion tasks compared to existing hardware benchmarks.
@@
Syntactically complete. Prior work (Ding et al., 2023) often selects random character spans as completion targets, which may break syntactic integrity. In contrast, recent studies (Liu et al., 2025b) show that developers expect LLMs to generate syntactically complete code units (e.g., full lines or blocks). MHRC-Bench-Eval uses tree-sitter [^1] https://github.com/tree-sitter/tree-sitter to parse each file into a concrete syntax tree and randomly selects a syntactic node as the completion target, using its exact source span as the ground-truth output.
@@
Data Independence and Inference difficulty. To ensure data independence and inference difficulty, we set the following rules: (1) The repositories and files used for the training, validation, and test sets are mutually exclusive, with no overlap. (2) We ensure that 40% of completion targets are between 2 and 5 lines inclusive. And test sets should be after 2022. (3) The target code should not exclusively include whitespace, debug message and comments. (4) For each file, we select exactly one target position. Therefore, we get 47,175 samples from 47,175 source files. The detailed information about data splits to make sure no overlap is shown in Appendix A.1.
@@ 3.4 Dataset Statistics
Table 2: Dataset statistics of MHRC-Bench. Val denotes validation.
Language Train Repos Val Repos Test Repos Train Files Val Files Test Files Cross-file Dependencies
Chisel 24 4 20 2,573 119 319 3.12
Verilog/SV 198 6 43 25,586 107 492 4.02
VHDL 173 7 59 13,039 104 491 4.34
HLS 22 3 25 3,766 123 456 5.85
@@ 3.5 Dataset Categories
To better analyze LLM’s performance in a fine-grained manner, we further provide two types of fine-grained annotations include code structural level and hardware semantic level for each completion target as shown in Figure 2.
@@
Hardware semantic level. We categorize the target code regions according to the hardware design and verification flow in Table 3, which broadly consists of design coding and testbench-based verification. Our semantic labels were generated through five hardware experts. Unlike software programs, where execution is inherently sequential, hardware code describes a parallel architecture or circuit with explicit structural, temporal, and interface semantics. This fundamental difference motivates a hardware-specific categorization of completion targets.
For design coding, we categorize target code into six classes: Design Structure, Declaration and Definition, Storage Block, Computation Block, Control Flow Block, and Interface, which capture key aspects of hardware design ranging from structural hierarchy and state storage to data-path computation, control logic, and communication interfaces. In addition, we explicitly model verification code which is critical in hardware development. Verification targets are further divided into Property and Assertion Specification, Testbench Stimulus and Environment Modeling, and Monitoring and Checking Logic, collectively characterizing how correctness properties are specified and how design behavior is observed and validated through simulation or formal verification. The distribution of hardware semantic categories is reported in Appendix A.3.
@@ 4.1 Evaluation Models and Metrics
Code model. We evaluate code LLMs of different scale. The evaluated models include DeepSeek-Coder-MQA-5.7B (Guo et al., 2024), DeepSeek-Coder-base-6.7B (Guo et al., 2024), Qwen2.5-Coder-3B-base (Hui et al., 2024), Qwen2.5-Coder-7B-base (Hui et al., 2024) and Qwen2.5-Coder-14B-base (Hui et al., 2024).
General model. We evaluate commercial models including GPT-5 (Singh et al., 2025), Gemini 2.5 Pro (Comanici et al., 2025), Grok4 (xAI, 2025), and Deepseek V3.2 (Liu et al., 2025a).
Metrics. We compare the generated code with the reference and compute the exact match (EM), and edit similarity (ES) following (Liu et al., 2025b; Zhang et al., 2023). Their detailed implementations are defined in A.4
@@ 4.2 Experimental Setup
Baseline and Tuning. We provide the model only with the file that contains the completion cursor. Without explicit inter-file context, the model must depend on its pretrained ability to produce the completion. To further improve the performance of repository-level code completion, we fine-tune code LLMs on training dataset supervised fine-tuning (SFT) with detailed configuration in Appendix A.4.
@@
Retrieval Framework and Prompt. There are many different repository-level code completion methods including RepoCoder (Zhang et al., 2023), GraphCoder (Liu et al., 2024c), and RLCoder (Wang et al., 2024). We evaluate these frameworks with different models on MHRC-Bench-Eval. All models use a fixed 2,048-token input window. The retriever-specific hyper parameters are systematically tuned per method according to their papers. We adopt two kinds of prompt, preceding-context and full-context (fill in the middle) shown in Appendix A.4, The preceding-context-only setting includes only preceding code and line of target code, while the full-context setting includes both preceding and succeeding code within the same window budget.
Table 4: EM % and ES % scores across four hardware description languages. Full context refers to the use of both preceding code and the succeeding code surrounding the target code segment, whereas preceding context-only uses only the preceding code as context.
Model Retrieval Method Preceding context Full context
Chisel V/SV VHDL HLS Chisel V/SV VHDL HLS
EM ES EM ES EM ES EM ES EM ES EM ES EM ES EM ES
GPT-5 No RAG 15.2 51.6 21.9 55.4 20.8 58.5 12.1 46.7 24.8 56.7 34.5 60.9 15.4 52.0 31.4 63.7
RepoCoder 18.2 54.1 22.5 56.1 22.4 60.3 14.8 47.2 27.6 59.6 35.4 61.3 16.7 54.1 34.4 63.9
GraphCoder 11.1 17.3 14.3 35.4 7.8 33.1 16.2 25.2 13.9 44.2 15.7 40.1 7.5 35.7 24.9 49.5
RLCoder 17.9 55.3 23.8 55.9 26.6 61.4 17.4 56.2 27.8 60.2 36.2 61.7 17.4 55.3 36.9 72.9
Gemini 2.5 Pro No RAG 11.1 56.2 28.9 61.4 23.3 61.8 18.0 52.0 33.2 59.9 33.2 59.9 16.3 50.8 34.6 63.8
RepoCoder 11.4 58.3 30.9 60.8 24.7 63.6 23.1 53.2 33.7 61.7 34.8 59.6 18.0 52.3 39.4 65.3
GraphCoder 2.5 17.2 2.6 26.5 3.1 25.8 5.3 17.7 13.3 50.8 22.2 50.5 9.1 45.0 29.1 57.8
RLCoder 12.8 59.2 30.7 62.0 25.9 64.3 24.2 55.4 34.6 63.2 35.3 60.2 19.2 53.7 41.2 66.9
Grok4 No RAG 10.2 44.3 15.6 49.0 10.4 46.1 9.0 40.3 22.4 53.7 31.4 56.8 14.3 47.8 31.6 60.5
RepoCoder 11.3 45.6 16.5 49.0 11.6 47.9 12.0 40.8 23.8 54.7 32.0 56.1 15.7 49.3 32.3 61.1
GraphCoder 2.2 24.1 6.4 23.4 0.6 24.3 6.4 23.4 12.7 35.4 16.5 38.1 8.1 35.3 18.0 38.8
RLCoder 13.7 46.5 17.3 50.7 13.2 48.4 12.4 42.4 25.6 55.2 33.4 57.2 16.8 51.2 33.3 62.2
Deepseek V3.2 No RAG 10.3 47.0 22.8 56.3 17.0 53.8 9.9 42.4 22.6 51.2 28.7 54.1 13.3 43.0 30.3 57.3
RepoCoder 10.6 48.3 23.5 57.2 17.6 54.6 11.2 43.3 23.2 52.1 28.3 52.6 14.2 42.8 33.3 57.9
GraphCoder 0.9 24.5 0.2 39.7 0.6 32.8 0.7 28.6 9.3 39.5 15.1 39.1 9.9 39.1 22.7 47.9
RLCoder 12.3 48.9 24.2 59.2 19.4 55.3 13.1 44.0 24.3 53.1 29.4 45.1 14.2 43.6 34.3 59.2
Table 5: EM (%) and ES (%) across different languages. Full context refers to the use of both preceding code and the succeeding code surrounding the target code segment, whereas preceding context-only uses only the preceding code as context.
Model Preceding context Full context
Chisel V/SV VHDL HLS Chisel V/SV VHDL HLS
EM ES EM ES EM ES EM ES EM ES EM ES EM ES EM ES
Qwen2.5-Coder-3B 0.0 15.5 0.0 19.0 0.0 22.8 0.2 15.7 0.0 7.9 0.2 7.9 0.0 6.6 0.0 5.5
+Tuning 17.3 56.1 28.9 61.5 26.6 64.7 17.1 54.0 24.8 54.3 34.8 65.5 28.2 59.1 33.1 55.0
+Retrieval & Tuning 19.0 57.6 30.2 63.4 28.3 66.2 18.3 56.0 26.3 56.1 36.2 67.0 29.6 60.8 34.7 56.9
Qwen2.5-Coder-7B 0.0 9.8 0.0 12.8 0.2 14.0 0.0 9.7 0.0 14.8 0.0 12.7 0.0 13.0 0.2 14.1
+Tuning 19.2 59.3 30.1 63.8 29.2 66.3 18.2 56.2 26.6 66.7 39.0 70.8 34.5 71.9 35.5 69.5
+Retrieval & Tuning 20.8 61.4 31.9 65.2 30.7 68.2 20.2 57.8 28.4 68.5 40.7 72.6 36.1 73.6 37.1 71.3
Qwen2.5-Coder-14B 0.0 15.5 0.0 16.1 0.0 23.3 0.2 10.4 0.0 16.5 1.0 18.5 0.4 15.8 1.1 16.0
+Tuning 22.0 61.1 30.1 65.3 31.5 68.8 20.8 57.8 31.3 68.9 41.9 74.4 35.3 73.8 38.2 71.5
+Retrieval & Tuning 24.0 62.7 31.7 67.4 33.0 70.0 22.4 59.8 33.1 70.6 43.8 76.1 37.0 75.4 40.0 73.2
DS-Coder-5.7B 0.0 22.9 0.4 24.5 0.4 30.1 0.2 16.9 0.0 13.2 0.4 8.8 0.0 9.0 0.0 9.7
+Tuning 16.7 54.6 28.7 63.2 27.6 65.9 17.9 54.9 24.8 61.2 37.6 71.7 30.7 71.1 33.1 67.7
+Retrieval & Tuning 18.5 56.2 29.9 65.3 29.5 67.3 19.5 56.9 26.5 63.0 39.2 73.5 32.4 72.8 34.8 69.3
DS-Coder-6.7B 0.31 20.7 0.0 21.5 0.2 28.3 0.0 15.5 0.0 5.4 0.2 5.7 0.0 5.6 0.0 3.6
+Tuning 14.6 52.4 28.5 63.7 26.6 65.2 20.0 55.7 25.7 56.5 37.8 67.4 31.5 61.1 36.8 60.4
+Retrieval & Tuning 16.5 53.9 30.0 65.6 28.6 67.0 21.5 57.6 27.3 58.4 39.5 69.2 33.1 63.0 38.6 62.1